OpenStudy (anonymous):

27x^2-3x+1=0 Someone please help me factor and solve for zero!

OpenStudy (ybarrap):

since (-3)^2 - 4*27*1 = 9 - 108 = -99 < 0, there are no real solutions.

OpenStudy (ybarrap):

See attached for solution

OpenStudy (anonymous):

thank you!